OpenStudy (anonymous):

The length of an arc defined by a 60o central angle in a circle of diameter 3 in. is ?

OpenStudy (anonymous):

length of complete circle with radius r: L= 2Pi*r 60º is 60/360=1/6 of a circle, so L =2Pi*3*1/6=Pi

OpenStudy (anonymous):

i dont understand ? @myko

OpenStudy (anonymous):

the formula for length of the circle is L= 2Pi*r. The full circle has 360º. So if you take only 60º it would be 1/6 of the circle. So 1/6 of it's length

OpenStudy (anonymous):

ohhh so the equations would b 2Pi*6?

OpenStudy (anonymous):

no.\[ 2\Pi*r \over 6\]

OpenStudy (anonymous):

whats the radius? isnt it just twice the diameter?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh, sry, i missed that in your question. Radius is 3/2 in this case

OpenStudy (anonymous):

so L=2Pi*3/6*2=Pi/2

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

yw